OpenStudy (anonymous):

12 of the 40 cars pulled by locomotive were tankers. A:what fraction of the cars were tankers. B: what percent of the cars were not tankers

OpenStudy (anonymous):

12/40 = 3/10 3/10 = 30% meaning 60% of the cars were not tankers.

OpenStudy (johnweldon1993):

Well...close...12 of 40 is indeed 3/10 which is 30% were tankers yes...but that would mean that 100% - 30% = 70% were NOT tankers...

OpenStudy (anonymous):

err 70 lmao

OpenStudy (anonymous):

@johnweldon1993 Thanks for catching that, bleh